Perfom the number operations on four 2s to make numbers 0 through 6.Use any of the following symbols; x,divide,+,-,().
Answer by Edwin McCravy(20086)


2+2-2-2 = 0
(2+2-2)÷2 = (4-2)÷2 = 2÷2 = 1
(2-2)÷2+2 = 0÷2+2 = 0+2 = 2
(2+2+2)÷2 = 6÷2 = 3
2+2+2-2 = 6-2 = 4
2÷2+2+2 = 1+2+2=5
2x2x2-2 = 8-2 = 6

Edwin